OpenStudy (anonymous):

6x^2+5=7x i have to use the quadratic formula to find the exact solutions... i keep trying with the formula and cant figure out the final answer!

OpenStudy (saifoo.khan):

use quadratic formuale/

OpenStudy (anonymous):

hey saifoo

OpenStudy (saifoo.khan):

\begin{array}l\color{red}{\text{h}}\color{orange}{\text{e}}\color{#9c9a2e}{\text{y}}\color{green}{\text{y}}\color{blue}{\text{y}}\color{purple}{\text{ }}\color{purple}{\text{S}}\color{red}{\text{u}}\color{orange}{\text{m}}\color{#9c9a2e}{\text{m}}\color{green}{\text{e}}\color{blue}{\text{r}}\color{purple}{\text{!}}\color{purple}{\text{}}\end{array}

OpenStudy (anonymous):

haha nice

OpenStudy (anonymous):

assume it is in the form of ax^2 + bx + c = 0 x1 = {-b + [sqrt(b^2 -4ac)]} / 2a x2 = {-b - [sqrt(b^2 -4ac)]} / 2a

OpenStudy (llort):

6x^2-7x+5=0 try it now

OpenStudy (anonymous):

so i use -(-7)+/-sqrt((-7)^2-4(6)(5)

OpenStudy (anonymous):

and divide it by 2(6) which is 12?

OpenStudy (anonymous):

yes